I was absent last week and its come back to bite me in the butt. This is the question best I know how to type it in. Mostly word for word.
If f(x)=3x-4 and g(x)=x+2 find f.g.
Thanks

Answer by jim_thompson5910(35256)

I'm assuming you want to multiply the equations

f*g = (3x-4)(x+2)

f*g = 3x^2 + 6x - 4x - 8

f*g = 3x^2 + 2x - 8